# Koopmon trajectories in nonadiabatic quantum-classical dynamics

### 21 Dec 2023 v1

@@ 3 Implementation and comparison to other methods
The koopmon scheme was implemented in MATLAB to probe its performance across various model systems. Both MATLAB versions 2021b and 2023b were used during the development of this work. In the following we provide a detailed summary of all the different input parameters:
@@ 3.1 Koopmon integration scheme
For the numerical integration of the koopmon equations presented in (14), we use a fourth-order Runge-Kutta method (RK4), for which we introduce the following input parameters:
@@
    Time increment: Depending on the timescales involved in the underlying model system, we use different time increments $dt>0$. The selected increments guarantee that the relative error of the time-evolved energy stays within the tolerance $10^{-2}$.
@@
    Initial conditions: The initial condition for the classical position of each trajectory in the koopmon method is sampled from a normal distribution $\mathcal{N}(\mu_{q},\sigma_{q}^{2})$, where the mean and the standard deviation $(\mu_{q},\sigma_{q})\in\mathbb{R}\times\mathbb{R}^{+}$ are given parameters. Analogously, the classical momenta are sampled from a normal distribution $\mathcal{N}(\mu_{p},\sigma_{p}^{2})$ for a given $\mu_{p}\in\mathbb{R}$ and $\sigma_{p}=1/(2\sigma_{q})$. In order to achieve fast convergence of the koopmon method with an increasing number $N$ of trajectories, we do not work with the internal MATLAB command for the generation of normally distributed random numbers. Instead, the sampling from the above normal distributions is performed using quasi-random numbers based on the two-dimensional Sobol sequence [50]. In the quantum sector, the initial condition for the density matrix was uniformly chosen as $\hat{\varrho}_{a}=\hat{\rho}_{0}$ for all $a=1,\dots,N$, where $\hat{\rho}_{0}\in\operatorname{Her}(\mathbb{C}^{2})$ is a given Hermitian matrix.
@@ 3.2 The quantum bohmion method
In order to assess the koopmon method in terms of its accuracy and computational costs, we will compare its results to those obtained by both the MQC Ehrenfest model and fully quantum descriptions. Indeed, coupled-trajectory particle schemes similar to the MQC koopmon method also appear in fully quantum approaches based on quantum hydrodynamics [3, 19]. In this context, one considers the coupled quantum dynamics of nuclei and electrons and describes the nuclear component by Madelung-Bohm trajectories, while treating the electrons in the standard Schrödinger picture. For example, the bohmion method [19, 36] is based on exactly the same equations as (14), although the total energy (15) is replaced by
@@ 3.3 Schrödinger solver and output visualization
As anticipated, we will analyze koopmon dynamics and compare its features to those appearing in SOFT, Ehrenfest, and bohmion dynamics. Various quantities are visualized in both the classical and quantum sectors. For the particle methods, we present the classical Liouville density by plotting the particle clouds $(\mathbf{q}(t),\mathbf{p}(t))\in\mathbb{R}^{N}\times\mathbb{R}^{N}$ at different times. In SOFT simulations, we plot the Wigner distribution of the time-evolved wavepacket $\Psi(t)$. Notice that, unlike the classical Liouville density, the quantum Wigner distribution can take negative values. As we will see, areas with a negative sign in the Wigner distribution are particularly crucial for analyzing the koopmon method. We recall that the negative values developed by Wigner dynamics are distinctive tracks of purely quantum features, typically related to interference effects. To enhance the visual comparison of the Wigner distribution with the classical particle clouds, the latter will be supplemented with the regularized density $D({\mathbf{z}},t):=(K^{(\Delta)}*f)({\mathbf{z}},t)=\sum_{a}w_{a}K^{(\Delta)}({\mathbf{z}}-{\boldsymbol{\zeta}}_{a}(t)),$ where $K^{(\Delta)}$ is a Gaussian kernel whose width $\Delta>0$ is a purely visualization parameter, independent of the model parameter $\alpha$. Here, we have fixed the value $\Delta=0.25$ for all test cases. In the quantum sector, we plot the evolution of the quantum purity $\|\hat{\rho}(t)\|^{2}=\operatorname{Tr}\hat{\rho}(t)^{2}$, which is a basis-independent measure of decoherence. Note that in SOFT simulations we have $\hat{\rho}(t)=\int\Psi(r,t)\Psi(r,t)^{\dagger}\,{\rm d}r$, while $\hat{\rho}(t)=\sum_{a}w_{a}\hat{\varrho}_{a}(t)$ for the particle methods.
@@ 4 Results for the Tully models
We start the presentation of our numerical results for the Tully models, originally introduced by John Tully to probe his fewest switches surface hopping algorithm [64]. Over the decades, these models have evolved into essential tools for testing novel MQC algorithms in computational chemistry [39]. Each Tully model serves as a platform for simulating different nonadiabatic molecular processes. For example, the bohmion scheme was successfully tested against all three Tully models, for which it exhibited great accuracy levels with a regularization parameter $\alpha\leq 1/20$. The first two models are designed to simulate the dynamics of a wavepacket entering a coupling region (I: single avoided crossing, II: double avoided crossing) on the lower surface. By passing through these coupling areas, parts of the wavepacket move from the lower to the upper surface. The third model is known to present a typically challenging test case, modeling not only the transmission of wavepackets as in Tully I and II, but also involving reflection effects. All results presented in this paper are based on initializations in low (I+II) or intermediate (III) momentum regimes, which are known to be more challenging test cases than those with high initial momentum.
@@
with model-dependent functions $H_{0}$, $H_{1}$, and $H_{3}$. Otherwise, except for the initial conditions, the same input parameters listed in Section 3 were used across all three Tully models. These are: Input 1) The classical subsystem is characterized by an effective mass $M=2000$. Input 2) The number of particles was set to $N=1000$. Input 3) The kernel width was set to $\alpha=0.325$. Input 6) The time increment was set to $dt=2$. Input 7) The width of the initial Gaussian wavepacket is linked to the initial momentum via the relation $\sigma_{q}=20/(\sqrt{2}\mu_{p})$ [3, 36]. The parameter values in items 2) and 6) allow us to run the simulations on a laptop machine, thereby avoiding the need for computer clusters. As for the modeling parameter $\alpha$, values as high as $\alpha=0.75$ were also used on some occasions and we found that $\alpha=0.5$ already allows for a good comparison with the fully quantum results in all the cases considered in this work. The same holds by taking the number of particles down to $N=500$. The final values $N=1000$ and $\alpha=0.325$, chosen to be the same for all Tully models, are the result of a compromise between the computational cost associated to the evaluation of the integrals ${\cal I}_{ab}$ in (13) and the need for satisfactory accuracy levels for all Tully models. Notice that the chosen value of $\alpha$ is well above the $\alpha=1/20$ used in previous bohmion simulations of the Tully models [36]. Indeed, in that case, such low values of $\alpha$ and their corresponding fine integration grid are required by the complexity of the integrals ${\mathscr{I}}_{ab}$ in (18), which comprise the regularization of the quantum potential.
@@ 4.1 Tully I: single avoided crossing
with $a=0.01,\,b=1.6,\,c=0.005,\,d=1$. A plot of the corresponding potential energy surfaces (PESs) can be found in the left box of Figure 1. These PESs are defined as the $q$-dependent eigenvalues of the $2\times 2$ matrix ${H_{0}(q){\boldsymbol{1}}+H_{1}(q)\hat{\sigma}_{x}+H_{3}(q)\hat{\sigma}_{z}}$. In the classical sector, the initial mean values for position and momentum (Input 7) are chosen as $(\mu_{q},\mu_{p})=(-8,10)$. Moreover, in the quantum sector, the initial density matrix represents the ground state, which corresponds to the choice $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(1,0)^{T}$. The same set of model parameters has been used previously by other authors [3, 36].
@@
Each row of Figure 2 shows a snapshot of the phase-space distribution obtained by a different method, up to the final time $t=3000$ of the simulation. As we can see for the Wigner distribution in the left column, the wavepacket starts from its initial position and evolves towards the right due to its positive initial momentum. When reaching the avoided crossing, nonadiabatic transitions occur, causing the wavepacket to split. Consequently, two regions with different momenta emerge, which are seen in the panel for the final time. We recall that energy crossings correspond to spectral degeneracies typically manifesting as points of intersection of the PESs. The regions where the PESs become close are referred to as avoided crossings.
Figure 2: Time evolution in the classical sector for Tully I. Each row shows a snapshot of the phase-space distribution. The first column displays the Wigner distribution of the fully quantum SOFT results, the second column refers to the koopmon dynamics of the phase-space density, while the third and fourth columns carry the analogous results for Ehrenfest and bohmion simulations, respectively.
For early simulation times, all particle methods are able to reproduce the shape of the Wigner distribution. Afterwards, the nonadiabatic transitions make it more challenging for all particle methods to accurately capture the dynamics. Also, we note that the Wigner distribution starts to develop negative values, identifying purely quantum effects. While Ehrenfest dynamics (third column) fail to reproduce the final configuration, we can see that for the bohmion method (last column) some particles have shifted towards regions of lower momentum. However, as hinted above, the chosen model parameter $\alpha=0.325$ is not small enough for bohmions to allow for the splitting to be captured precisely. The koopmon method (second column) yields the best results at all times. In particular, as we can see in the snapshot for the final time, a substantial number of particles have moved to the region of lower momentum and the onset of a splitting can be observed.
@@
The results in the quantum sector can be found in Figure 3. In addition to the purity plots at the right-hand side, we also illustrate the populations, which represent the proportion of the system in either the ground or the excited state. In the SOFT simulations, the population on the ground state is calculated as $P_{1}(t)=\int|\langle v_{1}(r)|\Psi(r,t)\rangle|^{2}\,\mathrm{d}r$, where $v_{1}(r)\in\mathbb{C}^{2}$ denotes the eigenvector of $\widehat{H}_{I}(r)$ corresponding to the smaller eigenvalue. For the particle methods, we have $P_{1}(t)=\sum_{a}w_{a}\langle v_{a,1}(t)|\hat{\varrho}_{a}(t)v_{a,1}(t)\rangle$, where $v_{a,1}(t):=v_{1}(q_{a}(t))$. The population on the excited state is given by $P_{2}(t)=1-P_{1}(t)$.
@@
The left-hand side population plots show that between times $t\approx 1000$ and $t\approx 2250$, the koopmon method (green) has difficulties to reproduce the values obtained by the quantum solution (red). We observe that in this time region the negative values of the Wigner function become distinctively visible in the fully quantum simulation: these negative values produce purely quantum effects that are not captured by the MQC koopmon method. However, for the final times, similar to the Ehrenfest (orange) and bohmion (blue) simulations, the koopmons successfully capture the correct population levels. The purity results show that the koopmon method yields good agreement with the quantum solution at all times. In particular, at the final time, we can see that the error for both the bohmions and the koopmons is very similar. However, the koopmons slightly overestimate, while the bohmions underestimate the correct purity levels. In comparison, Ehrenfest dynamics exhibit small overdecoherence, indicated by lower purity values.
@@ 4.2 Tully II: dual avoided crossing
with $a=0.05,\,b=0.28,\,c=0.015,\,d=0.06,\,e_{0}=0.025$. A plot of the corresponding PESs can be found in the middle panel of Figure 1. In the classical sector, the initial mean values for position and momentum (Input 7) are chosen as $(\mu_{q},\mu_{p})=(-8,16)$. Moreover, in the quantum sector, the initial density matrix represents the ground state, which corresponds to $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(1,0)^{T}$.
@@
Figure 4: Time evolution in the classical sector for Tully II. The presentation follows the same scheme as in Figure 2.
The second row shows that all particle methods yield similar results, accurately reproducing areas where the Wigner distribution attains positive values. Noticeable differences in the density of the particle clouds can be seen between Ehrenfest and bohmion dynamics, where particles are denser, and the koopmon method, for which the particles show a wider spread.
Similar to Tully I, both Ehrenfest and bohmion simulations fall short in reproducing the final wavepacket configuration. In contrast, for the koopmon method some of the particles are observed to move towards regions of lower momentum, resulting in a spreading of the particles rather than a definite splitting. We would like to note that further simulations indicate that increasing the number of particles in the koopmon method does not alleviate this effect. Indeed, the particle increase may need to be accompanied by a lower value of $\alpha$.

### 17 Jun 2024 v2

@@ 3 Implementation and comparison to other methods
The koopmon scheme was implemented in MATLAB to probe its performance across various model systems. Both MATLAB versions 2021b and 2023b were used during the development of this work. In what follows we provide a detailed summary of all the different input parameters:
@@ 3.1 Koopmon integration scheme
For the numerical integration of the koopmon equations presented in (14), we use a fourth-order Runge–Kutta method (RK4), for which we introduce the following input parameters:
@@
    Time increment: Depending on the model under consideration (Tully or Rabi), we work with different fixed time steps that ensure that the relative error of the time-evolved energy remains within the tolerance $10^{-2}$. The exact numerical values will be given later on in each case of study.
@@
    Initial conditions: The initial condition for the classical position of each trajectory in the koopmon method is sampled from a normal distribution $\mathcal{N}(\mu_{q},\sigma_{q}^{2})$, where the mean and the standard deviation $(\mu_{q},\sigma_{q})\in\mathbb{R}\times\mathbb{R}^{+}$ are given parameters. Analogously, the classical momenta are sampled from a normal distribution $\mathcal{N}(\mu_{p},\sigma_{p}^{2})$ for a given $\mu_{p}\in\mathbb{R}$ and $\sigma_{p}=1/(2\sigma_{q})$. In order to achieve fast convergence of the koopmon method with an increasing number $N$ of trajectories, we do not work with the internal MATLAB command for the generation of normally distributed random numbers. Instead, the sampling from the above normal distributions is performed using quasi-random numbers based on the two-dimensional Sobol sequence [58]. In the quantum sector, the initial condition for the density matrix was uniformly chosen as $\hat{\varrho}_{a}=\hat{\varrho}_{0}$ for all $a=1,\dots,N$, where $\hat{\varrho}_{0}\in\operatorname{Her}(\mathbb{C}^{2})$ is a given Hermitian matrix.
@@ 3.2 The quantum bohmion method
In order to assess the koopmon method in terms of its accuracy and computational costs, we will compare its results to those obtained by both the MQC Ehrenfest model and fully quantum descriptions. Indeed, coupled-trajectory particle schemes similar to the MQC koopmon method also appear in fully quantum approaches based on quantum hydrodynamics [3, 20]. In this context, one considers the coupled quantum dynamics of nuclei and electrons and describes the nuclear component by Madelung–Bohm trajectories, while treating the electrons in the standard Schrödinger picture. For example, the bohmion method [20, 38] is based on exactly the same equations as (14), although the total energy (15) is replaced by
@@ 3.3 Schrödinger solver and output visualization
As anticipated, we will analyze koopmon dynamics and compare its features to those appearing in SOFT, Ehrenfest, and bohmion dynamics. Various quantities are visualized in both the classical and quantum sectors. For the particle methods, we present the classical Liouville density by plotting the particle clouds $(\mathbf{q}(t),\mathbf{p}(t))\in\mathbb{R}^{N}\times\mathbb{R}^{N}$ at different times. In SOFT simulations, we plot the Wigner distribution of the time-evolved wavepacket $\Psi(t)$. More precisely, for the vector-valued wavefunction $\Psi=(\psi_{1},\psi_{2})^{T}$ we compute $W_{\Psi}:=W_{\psi_{1}}+W_{\psi_{2}}$, where $W_{\psi_{j}},\,j\in\{1,2\}$, denotes the usual Wigner distribution of the one-dimensional component $\psi_{j}\in L^{2}(\mathbb{R})$, given by
$W_{\psi_{j}}(q,p):=\frac{1}{\pi\hbar}\int_{-\infty}^{\infty}\psi_{j}^{*}(q+y)\psi_{j}(q-y)e^{2ipy/\hbar}\,\mathrm{d}y,\qquad\text{for all $(q,p)\in\mathbb{R}^{2}$.}$
In the reminder of this paper, the term wavepacket refers indifferently to either the wavefunction representation in configuration space or its Wigner distribution in phase space. Notice that, unlike the classical Liouville density, the quantum Wigner distribution can take negative values. As we will see, areas with a negative sign in the Wigner distribution are particularly crucial for analyzing the koopmon method. We recall that the negative values developed by Wigner dynamics are distinctive tracks of purely quantum features, typically related to interference effects. To enhance the visual comparison of the Wigner distribution with the classical particle clouds, the latter will be supplemented with the regularized density $D({\mathbf{z}},t):=(K^{(\Delta)}*f)({\mathbf{z}},t)=\sum_{a}w_{a}K^{(\Delta)}({\mathbf{z}}-{\boldsymbol{\zeta}}_{a}(t)),$ where $K^{(\Delta)}$ is a Gaussian kernel whose width $\Delta>0$ is a purely visualization parameter, independent of the model parameter $\alpha$. Here, we have fixed the value $\Delta=0.25$ for all test cases. In the quantum sector, we plot the evolution of the quantum purity $\|\hat{\rho}(t)\|^{2}=\operatorname{Tr}\hat{\rho}(t)^{2}$, which is a basis-independent measure of decoherence. Note that in SOFT simulations we have $\hat{\rho}(t)=\int\Psi(r,t)\Psi(r,t)^{\dagger}\,{\rm d}r$, while $\hat{\rho}(t)=\sum_{a}w_{a}\hat{\varrho}_{a}(t)$ for the particle methods.
@@ 4 Results for the Tully models
We start the presentation of our numerical results for the Tully models, originally introduced by John Tully to probe his fewest switches surface hopping algorithm [74]. Over the decades, these models have evolved into essential tools for testing novel MQC algorithms in computational chemistry [41, 57]. Each Tully model serves as a platform for simulating different nonadiabatic molecular processes; see Appendix A for further details from a chemical perspective. For example, the bohmion scheme was successfully tested against all three Tully models, for which it exhibited great accuracy levels with a regularization parameter $\alpha\leq 1/20$. The first two models are designed to simulate the dynamics of a wavepacket entering a coupling region (I: single avoided crossing, II: double avoided crossing) on the lower surface. By passing through these coupling areas, parts of the wavepacket move from the lower to the upper surface. The third model is known to present a typically challenging test case, modeling not only the transmission of wavepackets as in Tully I and II, but also involving reflection effects. All results presented in this paper are based on initializations in low (I+II) or intermediate (III) momentum regimes, which are known to be more challenging test cases than those with high initial momentum.
@@
with model-dependent functions $H_{0}$, $H_{1}$, and $H_{3}$. Otherwise, except for the initial conditions, the same input parameters listed in Section 3 were used across all three Tully models. These are: Input 1) The classical subsystem is characterized by an effective mass $M=2000$. Input 2) The number of particles was set to $N=1000$. Input 3) The kernel width was set to $\alpha=0.325$. Input 6) The time increment was set to $dt=2$. Input 7) The width of the initial Gaussian wavepacket is linked to the initial momentum via the relation $\sigma_{q}=20/(\sqrt{2}\mu_{p})$ [3, 38]. The parameter values in inputs 2) and 6) allow us to run the simulations on a laptop machine, thereby avoiding the need for computer clusters. As for the modeling parameter $\alpha$, values as high as $\alpha=0.75$ were also used on some occasions and we found that $\alpha=0.5$ already allows for a good comparison with the fully quantum results in all the cases considered in this work. The same holds by taking the number of particles down to $N=500$. The final values $N=1000$ and $\alpha=0.325$, chosen to be the same for all Tully models, are the result of a compromise between the computational cost associated to the evaluation of the integrals ${\cal I}_{ab}$ in (13) and the need for satisfactory accuracy levels for all Tully models. Notice that the chosen value of $\alpha$ is well above the $\alpha=1/20$ used in previous bohmion simulations of the Tully models [38]. Indeed, in that case, such low values of $\alpha$ and their corresponding fine integration grid are required by the complexity of the integrals ${\mathscr{I}}_{ab}$ in (18), which comprise the regularization of the quantum potential.
Figure 2: Time evolution in the classical sector for Tully I. Each column (1st: SOFT, 2nd: koopmons, 3rd: Ehrenfest, 4th: bohmions) shows a snapshot of the distribution in phase space. Momentum on the vertical axis, position on the horizontal axis (both in atomic units). The rows correspond to times $t=0$, $1280$, $2130$, and $3000$ (top to bottom, all in atomic units). In addition, $N=1000$ and $\alpha=0.325$.
@@ 4.1 Tully I: single avoided crossing
with $a=0.01,\,b=1.6,\,c=0.005,\,d=1$. A plot of the corresponding potential energy surfaces (PESs) can be found in the left box of Figure 1. These PESs are defined as the $q$-dependent eigenvalues of the $2\times 2$ matrix ${H_{0}(q){\boldsymbol{1}}+H_{1}(q)\hat{\sigma}_{x}+H_{3}(q)\hat{\sigma}_{z}}$. In the classical sector, the initial mean values for position and momentum (Input 7) are chosen as $(\mu_{q},\mu_{p})=(-8,10)$. Note that the chosen parameters result in the initial kinetic energy $E_{\operatorname{kin}}(0)=p_{0}^{2}/(2M)=0.025$, which lies above the upper PES. Moreover, in the quantum sector, the initial density matrix represents the ground state, which corresponds to the choice $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(1,0)^{T}$. The same set of model parameters has been used previously by other authors [3, 38].
@@
Each row of Figure 2 shows a snapshot of the phase-space distribution obtained by a different method, up to the final time $t=3000$ of the simulation. As described in Section 3.3, the particle clouds are overlapped to the convoluted density ${D:=K^{(\Delta)}*f}$ for visualization purposes. As we can see for the Wigner distribution in the left column, the wavepacket starts from its initial position and evolves towards the right due to its positive initial momentum. When reaching the avoided crossing, nonadiabatic transitions occur, causing the wavepacket to split. Consequently, two regions with different momenta emerge, which are seen in the panel for the final time. We recall that energy crossings correspond to spectral degeneracies typically manifesting as points of intersection of the PESs. The regions where the PESs become close are referred to as avoided crossings; see Figure 1.
For early simulation times, all particle methods are able to reproduce the shape of the Wigner distribution. Afterwards, the nonadiabatic transitions make it more challenging for all particle methods to accurately capture the dynamics. Also, we note that the Wigner distribution starts to develop negative values, identifying purely quantum effects. While Ehrenfest dynamics (third column) fail to reproduce the final configuration, we can see that for the bohmion method (last column) some particles have shifted towards regions of lower momentum. However, as hinted above, the chosen model parameter $\alpha=0.325$ is not small enough for bohmions to allow for the splitting to be captured precisely. The koopmon method (second column) yields the best results at all times. In particular, as we can see in the snapshot for the final time (last row), the number of particles that have moved into the region of lower momenta (below $p\approx 6$) is large enough to show the onset of a splitting, which, however, never becomes distinct. We have seen similar behaviour in other test cases where no significant mass transfer is observed in the exact quantum dynamics. Having said that, we also point out that other investigations within the intermediate momentum regimes show that, whenever the particles are distributed more equally across the two PES, the splitting is more distinct and the koopmons succeed in achieving more accurate results.
@@
The results in the quantum sector can be found in Figure 3. In addition to the purity plots at the right-hand side, we also illustrate the populations, which represent the proportion of the system in either the ground or the excited state. In the SOFT simulations, the population on the ground state is calculated as $P_{1}(t)=\int|\langle v_{1}(r)|\Psi(r,t)\rangle|^{2}\,\mathrm{d}r$, where $v_{1}(r)\in\mathbb{C}^{2}$ denotes the eigenvector of $\widehat{H}_{I}(r)$ corresponding to the smaller eigenvalue. For the particle methods, we have $P_{1}(t)=\sum_{a}w_{a}\langle v_{a,1}(t)|\hat{\varrho}_{a}(t)v_{a,1}(t)\rangle$, where $v_{a,1}(t):=v_{1}(q_{a}(t))$. In other words, the population of the ground state is given by the weighted sum of the individual population levels $P_{1,a}(t):=\langle v_{a,1}(t)|\hat{\varrho}_{a}(t)v_{a,1}(t)\rangle$ of each particle. The population on the excited state is given by $P_{2}(t)=1-P_{1}(t)$.
@@
The population plots on the left show that between times $t\approx 1000$ and $t\approx 2250$, the koopmon method (green) has difficulties to accurately reproduce the values obtained by the quantum solution (red). We observe that in this time region the negative values of the Wigner function become distinctively visible in the fully quantum simulation: these negative values produce purely quantum effects that are not captured by the MQC koopmon method. However, for the final times, similar to the Ehrenfest (orange) and bohmion (blue) simulations, the koopmons successfully capture the correct population levels. The purity results show that the koopmon method yields good agreement with the quantum solution at all times. In particular, at the final time, we can see that the error for both the bohmions and the koopmons is very similar. However, the koopmons slightly overestimate, while the bohmions underestimate the correct purity levels. In comparison, Ehrenfest dynamics exhibit small overdecoherence, indicated by lower purity values.
@@ 4.2 Tully II: dual avoided crossing
with $a=0.05,\,b=0.28,\,c=0.015,\,d=0.06,\,e_{0}=0.025$. A plot of the corresponding PESs can be found in the middle panel of Figure 1. In the classical sector, the initial mean values for position and momentum (Input 7) are chosen as $(\mu_{q},\mu_{p})=(-8,16)$. Note that the chosen parameters result in the initial kinetic energy $E_{\operatorname{kin}}(0)=0.064$, which again lies above the upper PES; see Figure 1. Moreover, in the quantum sector, the initial density matrix represents the ground state, which corresponds to $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(1,0)^{T}$.
@@
Figure 4: Time evolution in the classical sector for Tully II. Each column (1st: SOFT, 2nd: koopmons, 3rd: Ehrenfest, 4th: bohmions) shows a snapshot of the distribution in phase space. Momentum on the vertical axis, position on the horizontal axis (both in atomic units). The rows correspond to times $t=0$, $860$, $1140$, and $2000$ (top to bottom, all in atomic units). In addition, $N=1000$ and $\alpha=0.325$.
@@
The second row shows that all particle methods yield similar results, accurately reproducing areas where the Wigner distribution attains positive values. Noticeable differences in the density of the particle clouds can be seen between Ehrenfest or bohmion dynamics, where particles are denser, and the koopmon method, for which the particles show a wider spread.
Similar to Tully I, both Ehrenfest and bohmion simulations fall short in reproducing the final wavepacket configuration. In contrast, for the koopmon method some of the particles are observed to move towards regions of lower momentum, resulting in a spreading of the particles rather than a definite splitting. We would like to note that further simulations indicate that increasing the number of particles in the koopmon method does not alleviate this effect. Indeed, the particle increase may need to be accompanied by a lower value of $\alpha$. We emphasize that, similarly to what happens in Tully I, other investigations within the intermediate momentum regimes show that, whenever the particles are distributed more equally across the two PES, the splitting is more distinct and the koopmons succeed in achieving more accurate results.